13) (15 pts) Solve the given IVP. y + 2y + 2y = 10 sin(2t), y(0) = 1, y(0) = 0
Answer #1

Goren y + 2y + 2y = lo so(2k), 9(0)21 4(0)2 homogeneous Soulon y+2y + 2y =0 0+ 20+ 2.20 YLL); cēt cos (l) + czēt sinch) PaYpCk) = -2 Cos (2k) - I sin(24) 9(4) 2 cēt cosch) + Gé sinch) -2 co$(24) – Sonce+). 46)-1 I= .0; -2. se: c[-écoscey -esincPlease rate my answer